OpenStudy (noodls):

There are 12 different pizza toppings, and Arturo must rank his top 6 in order. How many different possible rankings are there?

OpenStudy (anonymous):

12 choices for number 1 11 for #2 10 for #3 9 for #4 8 for #5, and 7 for #6 by the counting principle there are \[12\times 11\times 10\times 9\times 8\times 7\] ways to rank them

OpenStudy (noodls):

it was that simple?? :) Thank you!!

OpenStudy (noodls):

is another way of doing this with following: P(n,r) n! / (n-r)!

OpenStudy (anonymous):

it is the same exactly that formula you wrote is just a formula, i wouldn't use it

OpenStudy (anonymous):

no one want to write \[\frac{12!}{(12-6)!}=\frac{12!}{6!}\] \[=\frac{12\times 11\times 10\times 9\times 8\times7\times 6\times 5\times 4\times 3\times 2\times 1}{6\times 5\times 4\times 3\times 2\times 1}\] \[=12\times 11\times 10\times 9\times 8\times 7\]

OpenStudy (anonymous):

the formula just tells you where to stop, don't get married to it, think instead